OpenStudy (anonymous):

Solve for n: 30/6 = n/7 A. n=35 B. n=34 C. n=36 D. n=37

OpenStudy (anonymous):

c

OpenStudy (anonymous):

and a medal

OpenStudy (natasha18):

The answer is actually a. Just plug the value in 30/6=35/7 5 = 5

OpenStudy (anonymous):

Ok thanks, I did get a after working it out again @Natasha18

OpenStudy (natasha18):

No problem :)